# On scalable and efficient training of diffusion samplers

### v2

@@ A.4 Task details
TorchANI potential for Alanine Dipeptide.
We leverage TorchANI [16], a PyTorch implementation of ANI deep-learning potentials trained on quantum-mechanical reference data, to accurately calculate molecular energies. It provides transferable machine learning potential trained on organic molecules for efficient energy and force evaluations with accuracy comparable to density-functional theory (DFT). In particular, TorchANI excels at modeling small to medium-sized organic molecules such as alanine dipeptide.
@@ B.1 Low-dimensional standard benchmarks
PIS+LP $-1.32\pm 0.07$ $2.42\pm 0.20$ $3.75\pm 0.22$ 300M $160.83\pm 0.41$ $180.49\pm 4.76$ $19.66\pm 4.78$ 300M
TB+LP $-0.35\pm 0.03$ $0.53\pm 0.04$ $0.87\pm 0.03$ 160M $161.42\pm 0.40$ $195.89\pm 8.14$ $34.37\pm 8.15$ 300M
FL-SubTB+LP $-0.36\pm 0.01$ $0.58\pm 0.08$ $0.94\pm 0.07$ 260M $160.74\pm 0.15$ $215.93\pm 4.52$ $55.19\pm 4.52$ 330M
TB+LS+LP $-0.38\pm 0.03$ $0.32\pm 0.02$ $0.69\pm 0.02$ 320M $162.95\pm 0.08$ $166.30\pm 0.11$ $3.35\pm 0.14$ 320M
TB+Expl+LP $-0.37\pm 0.01$ $0.32\pm 0.02$ $0.69\pm 0.02$ 300M $160.76\pm 0.13$ $215.92\pm 14.90$ $55.16\pm 14.90$ 300M
TB+Expl+LS+LP $-0.37\pm 0.01$ $0.34\pm 0.02$ $0.71\pm 0.02$ 320M $162.97\pm 0.06$ $166.25\pm 0.10$ $3.28\pm 0.12$ 320M
PIS $-2.03\pm 0.22$ $55.48\pm 10.71$ $57.50\pm 9.02$ 100M $159.71\pm 1.70$ $333.79\pm 3.98$ $174.08\pm 4.33$ 100M
TB $-1.35\pm 0.04$ $99.04\pm 6.01$ $100.40\pm 5.67$ 100M $160.58\pm 0.87$ $439.28\pm 166.52$ $278.70\pm 166.49$ 100M
TB+LS $-0.38\pm 0.03$ $0.83\pm 0.46$ $1.21\pm 0.38$ 290M $163.12\pm 0.10$ $166.05\pm 0.12$ $2.93\pm 0.16$ 290M
TB+Expl+LS $-0.38\pm 0.05$ $0.58\pm 0.34$ $0.96\pm 0.34$ 290M $160.87\pm 3.31$ $168.27\pm 1.49$ $7.40\pm 3.63$ 290M
@@
iDEM $-2.14\pm 0.45$ $12.75\pm 3.67$ $14.89\pm 3.70$ 300M $142.23\pm 0.40$ $211.56\pm 2.53$ $69.33\pm 2.56$ 300M
Ours (round 1) $-0.40\pm 0.01$ $0.33\pm 0.02$ $0.73\pm 0.02$ 6M $162.49\pm 0.05$ $166.60\pm 0.01$ $4.11\pm 0.05$ 9M
Ours (round 2) $-0.40\pm 0.03$ $0.33\pm 0.05$ $0.73\pm 0.05$ 12M $162.63\pm 0.01$ $166.48\pm 0.03$ $3.85\pm 0.03$ 20M
@@
We benchmark our framework on two standard low-dimensional tasks: 40GMM and Manywell-32. Consistent with the high-dimensional experiments, we report ELBO, EUBO, their gap, and the number of energy calls required during training. We employ the same baseline methods and trajectory configurations (including trajectory length and training objectives) as in the high-dimensional settings. We provide detailed configurations, including diffusion scales for each task, in Appendix A.
@@ Results.
As demonstrated in Table 6, our method achieves competitive performance on lower-dimensional standard tasks, producing EUBO and ELBO metrics comparable to the strongest baselines, while using significantly fewer energy calls. On the 40GMM task, despite some baselines reporting strong ELBO and EUBO scores, they notably fail to capture the mode located at the bottom-right corner (see Figure 5). In contrast, our framework reliably identifies all modes without sacrificing performance metrics. We report both the first-round and second-round performances of our method, showing that our method attains robust performance on low-dimensional tasks even in the first round, with a slight but consistent improvement observed in the second round.
@@ B.2 Debiasing of Learner from MCMC Searcher
To evaluate whether the Learner effectively debiases the samples collected by the Searcher, we compare kernel density estimations (KDE) of samples obtained by the AIS Searcher with those generated by the on/off-policy TB Learner on Manywell distributions. Figure 6 illustrates these KDE comparisons across dimensions 32, 64, and 128.

### v4

@@ A.5 Task details
TorchANI potential for peptides.
We leverage TorchANI [16], a PyTorch implementation of ANI deep‐learning potentials trained on quantum‐mechanical reference data, to accurately calculate molecular energies. It provides transferable machine learning potential trained on organic molecules for efficient energy and force evaluation with accuracy comparable to density‐functional theory (DFT). In particular, TorchANI excels at modeling small peptides.
@@ B.1 Low-dimensional standard benchmarks
PIS+LP $1.32\pm 0.07$ $2.42\pm 0.20$ $3.75\pm 0.22$ 300M $160.83\pm 0.41$ $180.49\pm 4.76$ $19.66\pm 4.78$ 300M
TB+LP $0.35\pm 0.03$ $0.53\pm 0.04$ $0.87\pm 0.03$ 160M $161.42\pm 0.40$ $195.89\pm 8.14$ $34.37\pm 8.15$ 300M
FL-SubTB+LP $0.36\pm 0.01$ $0.58\pm 0.08$ $0.94\pm 0.07$ 260M $160.74\pm 0.15$ $215.93\pm 4.52$ $55.19\pm 4.52$ 330M
TB+LS+LP $0.38\pm 0.03$ $0.32\pm 0.02$ $0.69\pm 0.02$ 320M $162.95\pm 0.08$ $166.30\pm 0.11$ $3.35\pm 0.14$ 320M
TB+Expl+LP $0.37\pm 0.01$ $0.32\pm 0.02$ $0.69\pm 0.02$ 300M $160.76\pm 0.13$ $215.92\pm 14.90$ $55.16\pm 14.90$ 300M
TB+Expl+LS+LP $0.37\pm 0.01$ $0.34\pm 0.02$ $0.71\pm 0.02$ 320M $162.97\pm 0.06$ $166.25\pm 0.10$ $3.28\pm 0.12$ 320M
PIS $2.03\pm 0.22$ $55.48\pm 10.71$ $57.50\pm 9.02$ 100M $159.71\pm 1.70$ $333.79\pm 3.98$ $174.08\pm 4.33$ 100M
TB $1.35\pm 0.04$ $99.04\pm 6.01$ $100.40\pm 5.67$ 100M $160.58\pm 0.87$ $439.28\pm 166.52$ $278.70\pm 166.49$ 100M
TB+LS $0.38\pm 0.03$ $0.83\pm 0.46$ $1.21\pm 0.38$ 290M $163.12\pm 0.10$ $166.05\pm 0.12$ $2.93\pm 0.16$ 290M
TB+Expl+LS $0.38\pm 0.05$ $0.58\pm 0.34$ $0.96\pm 0.34$ 290M $160.87\pm 3.31$ $168.27\pm 1.49$ $7.40\pm 3.63$ 290M
@@
iDEM $2.14\pm 0.45$ $12.75\pm 3.67$ $14.89\pm 3.70$ 300M $142.23\pm 0.40$ $211.56\pm 2.53$ $69.33\pm 2.56$ 300M
SGDS (round 1) $\mathbf{-0.40\pm 0.01}$ $\mathbf{0.33\pm 0.02}$ $\mathbf{0.73\pm 0.02}$ 6M $162.49\pm 0.05$ $166.60\pm 0.01$ $4.11\pm 0.05$ 9M
SGDS (round 2) $\mathbf{-0.40\pm 0.03}$ $\mathbf{0.33\pm 0.05}$ $\mathbf{0.73\pm 0.05}$ 12M $\mathbf{162.63\pm 0.01}$ $\mathbf{166.48\pm 0.03}$ $\mathbf{3.85\pm 0.03}$ 20M
@@
We benchmark our framework on two standard low-dimensional tasks: 40GMM and Manywell-32. Consistent with the high-dimensional experiments, we report ELBO, EUBO, their gap, and the number of energy calls required during training. We employ the same baseline methods and trajectory configurations (including trajectory length and training objectives) as in the high-dimensional settings. We provide detailed configurations, including diffusion scales for each task, in Table˜5.
@@ Results.
As demonstrated in Table˜6, our method achieves competitive performance on lower-dimensional standard tasks, producing EUBO and ELBO metrics comparable to the strongest baselines, while using significantly fewer energy calls. On the 40GMM task, despite some baselines reporting strong ELBO and EUBO scores, they notably fail to capture the mode located at the bottom-right corner (see Figure˜5). In contrast, our framework reliably identifies all modes without sacrificing performance metrics. We report both the first-round and second-round performances of our method, showing that our method attains robust performance on low-dimensional tasks even in the first round, with a slight but consistent improvement observed in the second round.
@@ B.2 Debiasing of Learner from MCMC Searcher
To evaluate whether the Learner effectively debiases the samples collected by the Searcher, we compare kernel density estimations (KDE) of samples obtained by the AIS Searcher with those generated by the on/off-policy TB Learner on Manywell distributions. Figure˜6 illustrates these KDE comparisons across dimensions 32, 64, and 128.
@@
B.3 RND Weight Calibration
To evaluate the sensitivity of our method to the choice of the RND weight, we conducted additional experiments across approximately four different RND weight values per task. The results, summarized in Table Table˜7, show that the performance is largely robust to this hyperparameter. The calibration of the RND weight is straightforward: we run $20\%$ of the second-round training to estimate performance and select a reasonable value. As shown in the ELBO–EUBO gap measured at this stage, reported for Manywell-128, LJ potentials, and peptides, the variation across different RND weights is minor, indicating that extensive hyperparameter tuning is unnecessary.
Overall, these results suggest that the method performs consistently across a wide range of RND weights, with negligible degradation in stability or performance. Although a more systematic tuning strategy could be explored in future work, the current approach provides reliable results with minimal calibration effort.
Table 7: EUBO–ELBO gap at $20\%$ of the second-round training for different RND weight values.
RND weight $10^{0}$ $10^{1}$ $10^{2}$ $10^{3}$
Manywell-128 549.54 542.60 $\mathbf{538.79}$ 570.49
LJ-13 3.12 $\mathbf{2.67}$ 3.90 4.77
LJ-55 $\mathbf{34.76}$ 36.72 44.78 51.03
ALDP 17,531.95 $\mathbf{17,381.04}$ 17,417.75 17,410.98
B.4 Consistency with parallel tempering MD
We further validate the consistency of our method by replacing the Searcher with parallel tempering MD [40]. Parallel tempering (or replica exchange) MD is an enhanced sampling method that runs multiple independent simulations (or replicas) at different temperatures and periodically attempts to exchange them, allowing low-temperature simulations to overcome energy barriers and escape local minima. Also, we align the number of energy evaluations by collecting additional data for MLE, ensuring a fair comparison. Table˜8 shows the ELBO, EUBO, and their gap, demonstrating that our method achieves better metrics than MLE. This result indicates that our approach consistently works even when combined with advanced sampling techniques such as parallel tempering.
Table 8: Comparison of main metrics between parallel tempering + MLE (forward KL) training and our method with parallel tempering Searcher. We align the number of energy calls by collecting more data for MLE.
Method ELBO $\times 10^{3}$] $\uparrow$ EUBO $\times 10^{3}$] $\downarrow$ Gap $\times 10^{3}$] $\downarrow$
MLE $520.71\pm 0.02$ $538.03\pm 0.00$ $17.32\pm 0.02$
SGDS $521.03\pm 0.01$ $538.02\pm 0.00$ $16.99\pm 0.01$